# A Detailed Factor Analysis for the Political Compass Test: Navigating Ideologies of Large Language Models

### v1

@@ Abstract
Political Compass Test (PCT) or similar questionnaires have been used to quantify LLM’s political leanings. Building on a recent line of work that examines the validity of PCT tests, we demonstrate that variation in standard generation parameters does not significantly impact the models’ PCT scores. However, external factors such as prompt variations and fine-tuning individually and in combination affect the same. Finally, we demonstrate that when models are fine-tuned on text datasets with higher political content than others, the PCT scores are not differentially affected. This calls for a thorough investigation into the validity of PCT and similar tests, as well as the mechanism by which political leanings are encoded in LLMs.
@@ 1 Introduction
A large number of recent studies Feng et al. (2023); Motoki et al. (2024); He et al. (2024) use the Political Compass Test [^1] https://www.politicalcompass.org or PCT, a collection of 62 multiple-choice questions, where the respondent must agree on a Likert Scale (strongly disagree to strongly agree). These responses are then aggregated  [^2] The aggregation function is not public. to generate two distinct scores, a social score and an economic score, each ranging from $-10$ to $+10$. LLMs are generally prompted with each statement (possibly phrased as a question), and their level of agreement is recorded to infer the ideological coordinates (Figure 1).
@@
Figure 1: PCT scores in Mistral-7B-Instruct-v0.3 model before and after finetuning with multiple datasets with various generation parameters and prompts. We systematically investigate the effect of these factors on these scores.
@@
Theoretical validity issues Faulborn et al. (2025) aside, PCT suffers from empirical instability when used with LLMs. For example, Röttger et al. (2024) shows that the models’ answers flip when they are forced into the PCT’s multiple-choice format and change again with minimal paraphrases or in open-ended settings, revealing large prompt-sensitivity and low test–retest reliability. However, despite these criticisms, PCT is still used in recent papers Liu et al. (2025); Ye et al. (2025), and few studies have systematically evaluated the internal and external factors that can affect an LLM’s text generation, and consequently affect its PCT score. We bridge this gap by investigating two research questions:
@@
Which common decoding parameters, if any, affect PCT results? Decoding parameters do have a substantial effect on generations, but how that translates to the final PCT results is underexplored. We use one-way ANOVA tests on five common LLMs with varying sizes and four standard decoding parameters and find that the number of beams significantly affects the PCT results for some of the models, but overall, these parameters don’t affect the scores much. However, the prompt variation, as expected Röttger et al. (2024), has very strong effects (§3).
@@
How does fine-tuning affect PCT? This research question has two motivations. On the operational side, the parameter changes induced by fine-tuning naturally alter a model’s generation, but how that affects the PCT scores is unknown. Fine-tuning should not have any effect when controlled for prompt variations, as it introduces little information that can alter a model’s political leanings. However, we do find evidence of significant effects. On the cognitive side, this raises a question of whether this could be attributed to the text on which the models are fine-tuned. Specifically, we use two types of fine-tuning datasets – ones that contain political text, and ones that don’t. Arguably, human political leanings can change if new information is presented, and we hypothesize that fine-tuning serves as a good proxy for the same process in the models. We create a large collection of $\approx 2K$ PCT tests by fine-tuning multiple LLMs on eight datasets, but can not find a significant effect of the dataset type (§4).
@@
Beyond the general concern about the validity of PCT, our conclusions are: a) The LLM PCT scores are possibly robust against variation in the generation parameters, and b) Studies using PCT and other similar tests should verify their conclusions against both base and fine-tuned models. We hope this study inspires further investigation into how fine-tuning changes the political leanings of LLMs.
@@ 2 Experiment Setup
We use five open-source LLMs: Llama3-8B-Instruct Grattafiori et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), Falcon3-7B-Instruct Almazrouei et al. (2023), phi-4 Abdin et al. (2024), Gemma-3-4b-it Team et al. (2025). These models are widely used for chat and instruction-based applications and are well-known for their instruction-following capabilities. [^3] We use the smaller versions of these models as we fine-tune them later, but previous work has not found the scale to be a determining factor for PCT scores either Röttger et al. (2024). For all experiments, we prompt (eg. “Choose one of the following options”) the models with the PCT test statements (eg., “I’d always support my country, whether it was right or wrong”) and generate responses that we post-process and send to the PCT server, and get back the scores (see the appendix for details).
3 RQ1: Decoding Parameters & Prompting
@@
Our first experiment is to investigate the effect of standard decoding parameters on the PCT tests. We use the ten prompts described in Röttger et al. (2024), and for each prompt, we generate responses from the models by varying the following decoding parameters: top_k, temperature, num_beams, and num_beam_groups. top_k constrains the decoding probability space to the most important k tokens. A higher temperature value increases the variability of generation. A higher number of beams improves the quality at the possible cost of diversity. The num_beam_groups parameter determines the number of groups into which the beams would be divided – ensuring better diversity even in the case of higher num_beams. We choose $2-3$ values for each parameter, resulting in a total of $820$ PCT values across the five models.
@@
We assume these factors (and the prompts) should not have interaction effects (eg., the number of beams should not depend on the prompts or vice versa); therefore, we run one-way ANOVA tests using the social scores and economic scores as dependent variables and the decoding parameters as the independent ones. We use Levene’s test Levene (1960) to determine if the group variances are equal, and use Welch’s one-way ANOVA test Welch (1951) (which re-normalizes the degrees of freedom) when they are not.
Social Economic
Decoding Param Model F-statistic p-value F-statistic p-value
temp Gemma 1.5e-1 8.6e-1 8.9e-2 9.2e-1
Llama3 6.7e-1 5.1e-1 7.9e-1 4.6e-1
Falcon 5.0e-1 6.1e-1 1.9e-2 9.8e-1
Mistral 2.6e+0 7.5e-2 1.1e+0 3.5e-1
Phi 2.9e-5 1.0e+0 1.3e-3 1.0e+0
top k Gemma 2.3e-3 9.6e-1 4.3e-4 9.8e-1
Llama3 8.4e-2 7.7e-1 9.3e-2 7.6e-1
Falcon 2.0e-2 8.9e-1 4.3e-2 8.4e-1
Mistral 3.9e-2 8.4e-1 4.5e-3 9.5e-1
Phi 2.9e-5 1.0e+0 1.3e-3 9.7e-1
num Gemma 6.5e-1 5.2e-1 1.4e-1 8.7e-1
beams Llama3 1.2e+0 3.0e-1 1.2e+1 2.0e-5
Falcon 5.0e+1 6e-16 6.7e+1 5e-22
Mistral 2.8e+0 6.6e-2 2.8e-1 7.6e-1
Phi 1.3e-1 8.8e-1 1.5e-1 8.6e-1
num Gemma 2.0e-1 6.5e-1 7.7e-2 7.8e-1
beam Llama3 2.2e+0 1.4e-1 1.2e+1 5.7e-4
groups Falcon 1.2e-1 7.3e-1 1.4e+1 2.3e-4
Mistral 2.3e+0 1.3e-1 2.5e-1 6.2e-1
Phi 2.5e-2 8.7e-1 8.6e-3 9.3e-1
Table 1: One-way ANOVA factor analysis for generation parameters – bold denotes significant ones (p-value $<0.05$).
The results are presented in Table 1. Most parameters don’t have significant effects consistently across all models and scores, except for num_beams, which has a significant impact in Falcon (p-value $<0.05$).
Prompting has been shown to change the PCT responses Röttger et al. (2024), and our analysis indeed confirms this, as for all models, the factor has a significantly low p-value in both social and economic scores as well as a high F-statistic (detailed results in the appendix).
@@ 4 RQ2: Fine-Tuning
Having established that the decoding parameters don’t have a significant effect on the PCT tests, our next goal is to analyze the impact of fine-tuning. We investigate a diverse set of four natural language processing tasks (a) Classification, b) Conversation, c) Question-Answering, and d) Summarization) and eight distinct datasets for fine-tuning. For each of these tasks, we fine-tune the models with a control and a target dataset. A control dataset has textual content that is supposed to be neutral, i.e., non-politically oriented, so it should not impact the PCT scores. The target datasets, on the other hand, have text with strong political connotations, which could affect the trained models’ PCT score.
Model Social Economic
t-statistic p-value t-statistic p-value
Gemma -7.96 1.65e-14 8.65 8.28e-17
Llama3 -2.81 0.00527 -2.44 0.0153
Falcon 4.66 3.74e-06 -2.32 0.0211
Mistral -5.99 6.25e-09 -7.19 2.18e-12
Phi 8.74 3.48e-17 5.06 8.59e-07
@@
Gemma 40 4.4e-59 65 2.6e-15 5 4.0e-6 25 5.4e-38 101 1.5e-22 12 2.2e-17
Llama3 220 3.3e-209 47 1.7e-11 2 1.2e-2 18 8.0e-27 8 4.5e-3 5 2.9e-6
Falcon 12 1.8e-18 0.19 6.7e-1 1 4.6e-1 12 1.1e-17 10 1.4e-3 1 3.0e-1
Mistral 110 2.0e-131 131 2.8e-28 7 2.3e-10 40 1.8e-58 9 2.3e-3 4 2.0e-4
Phi 19 3.6e-27 50 7.0e-12 2 4.0e-2 39 1.1e-51 69 1.2e-15 12 2.4e-14
@@
Table 3: Two-way ANOVA results showing effects of prompt & finetuning (& their interaction) on Social and Economic axes across different models with non-significant effects italicized. F-statistics are rounded to save space.
@@
The details of the training process are described in the appendix. In essence, we produce nine model instances for each model class (Llama3/Phi etc.) – one is the base model, and the other eight are its fine-tuned versions on the eight datasets. We do not produce multiple model instances for the same base model and fine-tuning dataset by varying the initialization process, as our experiments suggest they are functionally equivalent. [^4] We train three instances of each model class on the SciSumm dataset using different seeds, but their test results do not vary significantly, see the appendix. We produce the PCT scores for these models by varying the prompts and other parameters as before, yielding a total of $3660$ PCT test results across the base and the fine-tuned models.
@@
First, we want to understand if the process of fine-tuning itself has an effect on PCT scores. We find that to be true – the average PCT scores on the social and economic axes differ significantly across the base vs fine-tuned versions of the models as measured by independent t-tests Virtanen et al. (2020). Specifically, Table 2 shows that all differences are statistically significant (p-value < $0.05$).
@@
However, it is expected that the PCT score of the fine-tuned model will depend on the prompt, and we are interested in observing the effect of fine-tuning while considering the effect of prompts. Therefore, we use two-way ANOVA tests with two independent variables: a) a categorical variable recording the prompt variation, and b) a binary variable indicating whether the model was fine-tuned or not. We test for the homogeneity of variances (Levene’s test) and normality of residuals (Shapiro–Wilk test Shapiro and Wilk (1965)), and when these conditions are violated, we use the Aligned Rank Transformed (ART) ANOVA Wobbrock et al. (2011) that first adjusts (or aligns) the data, then applies average ranks, allowing standard ANOVA methods to be used afterward.
@@
diff p-value diff p-value
Gemma 1.10e+01 0.1e-15 -1.00e+01 0.1e-15
Llama3 6.81e-02 2.63e-01 -9.05e-02 2.41e-01
Falcon 2.32e-02 7.70e-01 2.44e-01 5.31e-03
Mistral -1.46e-01 9.55e-02 2.45e-01 6.81e-02
Phi -6.47e-02 2.76e-01 3.44e-01 3.61e-03
Table 4: Games-Howell Test for mean difference in PCT results across control and target dataset groups.
Table 3 shows the results. Individually, both prompting and fine-tuning have significant effects, as does their interaction. In many cases (Llama3, Mistral), prompting has a more substantial impact than fine-tuning. Importantly, fine-tuning should not have any effect on a model’s political leanings, but that happens to be the case. We conclude that studies examining the validity of PCT and similar tests should also consider the scores from fine-tuned versions of the base models.
We investigate whether the dataset on which a model is fine-tuned contributes to the PCT score change by comparing the group mean differences of control versus target datasets using the Games-Howell test Games and Howell (1976), which accounts for heteroscedasticity in our data. Table 4 shows that we can only find a significant difference in one model (Gemma). If we consider PCT to be a valid test, this necessitates further investigation into the mechanism by which fine-tuning alters the political leanings of LLMs.

### v2

@@ Abstract
Political Compass Test (PCT) or similar questionnaires have long been used to quantify LLM’s political leanings. Building on a recent line of work that examines the validity of PCT tests, we demonstrate that variation in standard generation parameters does not significantly impact a model’s PCT scores. However, external factors such as prompt variations and fine-tuning individually and in combination affect the same. Finally, we demonstrate that when models are fine-tuned on text datasets with higher political content than others, the PCT scores are not differentially affected. We also generalize these findings to a similar popular test. This calls for a thorough investigation into the validity of PCT and similar tests, as well as the mechanism by which political leanings are encoded in LLMs.
@@ 1 Introduction
A large number of recent studies Feng et al. (2023); Motoki et al. (2024); He et al. (2024) use the Political Compass Test [^1] https://www.politicalcompass.org or PCT, a collection of 62 multiple-choice questions, where the respondent must agree on a Likert Scale (strongly disagree to strongly agree). These responses are then aggregated  [^2] The aggregation function is not public. to generate two distinct scores, a social score and an economic score, each ranging from 10-10- 10 to +10+10+ 10. LLMs are generally prompted with each statement (possibly phrased as a question), and their level of agreement is recorded to infer the ideological coordinates (Figure 1).
@@
Figure 1: Example PCT scores in Mistral-7B-Instruct-v0.3 model before and after finetuning with multiple datasets with various generation parameters but the same prompt(prompt 9)A.8. We systematically investigate the effect of these factors on these scores.
@@
Theoretical validity issues Faulborn et al. (2025) aside, PCT has been shown to suffer from empirical instability when used with LLMs. For example, Röttger et al. (2024) shows that the models’ answers flip when they are forced into the PCT’s multiple-choice format and change again with minimal paraphrases or in open-ended settings, revealing large prompt-sensitivity and low test–retest reliability. This problem has also been discussed in the context of framing bias Bang et al. (2024), where the authors proposed an alternate mechanism to understand models’ political bias. However, despite these criticisms, PCT is still used in recent papers Liu et al. (2025); Ye et al. (2025); Rozado (2024), and few studies have rigorously evaluated the internal and external factors that can affect an LLM’s text generation, and consequently, affect its PCT score. We bridge this gap by investigating two research questions:
@@
Which common decoding parameters, if any, affect PCT results? Decoding parameters do have a substantial effect on generations, but how that translates to the final PCT results is underexplored. One can illustrate how the scores change (Figure 1), but are these differences significant and systematic? We answer using ANOVA tests on five common LLMs with varying sizes and three standard decoding parameters and find that the number of beams significantly affects the PCT results for some of the models, but overall, these parameters don’t affect the scores much. However, the prompt variation, as expected Röttger et al. (2024), has strong effects (§3).
@@
How does fine-tuning affect PCT? This research question has two motivations. On the operational side, the parameter changes induced by fine-tuning naturally alter a model’s generation, but how that affects the PCT scores is unknown. Fine-tuning should not have any effect when controlled for prompt variations, as it introduces little information that can alter a model’s political leanings. However, we do find evidence of significant effects (for illustrative purposes, see Figures 2 and 3 in the appendix A.2). On the cognitive side, this raises a question of whether this could be attributed to the text on which the models are fine-tuned. Specifically, we use two types of fine-tuning datasets – ones that contain political text, and ones that don’t. Arguably, human political leanings can change in response to new information, and we hypothesize that fine-tuning serves as a good proxy for this process in the models. We create a large collection of $\approx 3K$ PCT tests by fine-tuning multiple LLMs on eight datasets, but can not find a significant effect of the dataset type (§4).
PCT is one among many benchmarks for measuring political leanings in humans that have been studied in the context of LLMs Rozado (2024). We reproduce our findings on “8 Values Test” IDRlabs (2023), another such popular benchmark, highlighting the generalizability of our work. Similar to the PCT, the 8 Values Test also degenerates into LLMs/humans producing four scores across four axes (by answering a set of questions, A.4), that we use as dependent variables in our analyses.
@@
Beyond the general concern about the validity of PCT and similar tests, our conclusions are: a) The LLM PCT scores are possibly robust against variation in the generation parameters, and b) Researchers should verify their conclusions against both base and fine-tuned models. We hope this study inspires further investigation into how fine-tuning changes the political leanings of LLMs.
@@ 2 Experiment Setup
We use four open-source LLMs: Llama3-8B-Instruct Grattafiori et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), Falcon3-7B-Instruct Almazrouei et al. (2023), and Gemma-3-4b-it Team et al. (2025). These models are widely used for chat and instruction-based applications and are well-known for their instruction-following capabilities.  [^3] We use the smaller versions of these models as we fine-tune them later, but previous work has not found the scale to be a determining factor for PCT scores either Röttger et al. (2024). Also, we use 4-bit quantized versions of these models. For all experiments, we prompt (eg. “Choose one of the following options”) the models with the PCT/8 Values test statements (eg., “I’d always support my country, whether it was right or wrong”) and generate responses that we post-process and send to the PCT/8 Values server, and get back the scores.
3 RQ1: Decoding Params & Prompting
@@
Our first experiment is to investigate the effect of standard decoding parameters on the PCT/8 Values tests. We use the ten prompts described in Röttger et al. (2024), and for each prompt, we generate responses from the models by varying the following decoding parameters: top_k, temperature, and num_beams. top_k constrains the decoding probability space to the most important k tokens. A higher temperature value increases the variability of generation. A higher number of beams improves the quality at the possible cost of diversity. We choose 222 values for each parameter.
@@
We assume that these factors (and the prompts) should not have interaction effects (eg., the number of beams should not depend on the prompts or vice versa); therefore, we run one-way ANOVA tests using the social scores and economic scores as dependent variables (8 Values have equivalent variables, A.4) and the decoding parameters as the independent ones. We use Levene’s test Levene (1960) to determine if the group variances are equal, and use Welch’s one-way ANOVA test Welch (1951) (which re-normalizes the degrees of freedom) when they are not.
The results are presented in Tables 3 (PCT, A.3) and 7 (8 Values, A.4). For 8 Values, none of the parameters has a significant effect for any model, and for PCT, only num_beams has a significant impact in Falcon (p-value <0.05<0.05< 0.05).
Confirming prior work’s conclusions Röttger et al. (2024), we find that prompting has a significantly low p-value (i.e., strong effect) in Economic scores as well as a high F-statistic (Table 4, A.3). However, we do not see such significance across the board for Social values. For 8 Values (Table 8, A.4), however, the prompts significantly affect all dependent variables across all models.
@@ 4 RQ2: Fine-Tuning
Having established that the decoding parameters don’t have a significant effect on the PCT tests, our next goal is to analyze the impact of fine-tuning. We investigate a diverse set of four natural language processing tasks (a) Classification, b) Conversation, c) Question-Answering, and d) Summarization) and eight distinct datasets for fine-tuning. For each of these tasks, we fine-tune the models with a control and a target dataset. A control dataset has textual content that is supposed to be neutral, i.e., non-politically oriented, so it should not impact the PCT scores. The target datasets, on the other hand, have text with strong political connotations, which could affect the trained models’ PCT score. The details of the datasets used in the experiments are provided in A.1.
@@
Gemma 29.1 4.15e-43 3.28e+01 1.55e-08 2.21 1.97e-02 21.5 2.77e-32 44.1 6.29e-11 4.57 6.99e-6
Llama3 4.51 8.75e-06 2.73e+01 2.32e-07 9.61e-01 4.72e-1 4.88 2.39e-06 44 6.62e-11 1.89 5.07e-2
Falcon 9.20 3.89e-13 1.90e+01 1.56e-05 4.85e-01 8.85e-01 8.30 1.02e-11 2.28e-02 8.80e-1 1.13 3.38e-01
Mistral 3.91 6.00e-05 1.33e+01 2.88e-04 3.42 4.47e-04 11.2 2.68e-16 221 6.75e-43 2.91 2.23e-3
@@
Table 1: Two-way ANOVA results showing effects of prompt & finetuning (& their interaction) on Social and Economic axes across different models with significant effects italicized. F-statistics are rounded to save space.
@@
The details of the training process are described in the appendix A.6, and the evaluation results are discussed in A.7. In essence, we utilize PEFT methods that modify the parameters of attention matrices and generate nine model instances for each model class (Llama3/Mistral, etc.). One instance is the base model, and the other eight are its fine-tuned versions on the eight datasets. [^4] We do not produce multiple model instances for the same base model and fine-tuning dataset by varying the initialization process, as our experiments suggest they are functionally equivalent. We train three instances of each model class on the SciSumm dataset using different seeds, but their test results do not vary significantly as illustrated in Table 18. We produce the PCT scores for these models by varying the prompts and other parameters as before, yielding a total of 269326932693 PCT test results across the base and the fine-tuned models.  [^5] Ideally, we should generate 288028802880 results (444 model types, 999 models of each type, 888 combinations of decoding params, and 101010 prompts), but some finetuned models are unable to generate responses for some PCT questions, and hence, are discarded. This number is 270427042704 for 8 Values.
@@
First, we aim to determine if the process of fine-tuning itself has an impact on PCT/8 Values scores. We find that to be true – the average PCT scores on the social and economic axes (and for equivalent variables in 8 Values) differ significantly across the base vs fine-tuned versions of the models as measured by independent t-tests Virtanen et al. (2020). See Table 5, A.3 for PCT and Table 9, A.4 for 8 Values.
@@
However, it is expected that the PCT/8 Values score of the fine-tuned model will depend on the prompt, and we are interested in observing the effect of fine-tuning while considering the effect of prompts. Therefore, we use two-way ANOVA tests with two independent variables: a) a categorical variable recording the prompt variation, and b) a binary variable indicating whether the model was fine-tuned or not. We test for the homogeneity of variances (Levene’s test) and normality of residuals (Shapiro–Wilk test Shapiro and Wilk (1965)), and when these conditions are violated, we use the Aligned Rank Transformed (ART) ANOVA Wobbrock et al. (2011) that first adjusts (or aligns) the data, then applies average ranks, allowing standard ANOVA methods to be used afterward.
Table 1 and Table 10 (A.4) show the results for PCT and 8 Values, respectively. Individually, both prompting and fine-tuning have significant effects, as does their interaction. Importantly, fine-tuning should not have any effect on a model’s political leanings, but that happens to be the case. We conclude that studies examining the validity of PCT and similar tests should also consider the scores from fine-tuned versions of the base models.
We hypothesize that changes in PCT/8 Values scores stem from the finetuning data: control-trained models should align with the base model, while target-trained ones should differ. To test this, for each task, we compute the group mean differences between the PCT/8 Values scores for base models and models trained on target or control datasets using the Games-Howell test Games and Howell (1976), which accounts for heteroscedasticity in our data. The results are presented in Tables 6 (PCT) and 12 (8 Values). For example, for the classification task, Gemma fine-tuned on the control dataset shows only a marginal difference from the base model on the PCT Social Score ($\text{p}=4.19\mathrm{e}{-2}$), while fine-tuning on the target dataset yields a highly significant shift ($\text{p}=3.53\mathrm{e}{-14}$).
Tables 2 (PCT) and 11 (8 Values, A.4), derived from Tables 6 and 12, show the fraction of tasks where the finetuned model exhibits such a significant shift from the base model. Surprisingly, we observe that the distinction in content does not matter, i.e., the models change their scores independently of the content they are fine-tuned on. This creates an opportunity for exploring the mechanism by which finetuning changes the political leaning of LLMs, as measured by PCT and similar tests, which we leave for future work.
@@
control target control target
Gemma 75% 75% 75% 75%
Llama3 100% 75% 75% 50%
Falcon 100% 75% 25% 50%
Mistral 100% 100% 100% 100%
Table 2: The fraction of tasks where finetuning significantly changes the PCT score of the models.
Effect of model size & quantization. Given the computational cost of finetuning, we use one model size per family and its 4-bit quantized version. A natural question is whether the findings can be generalized to larger models and their non-quantized versions. To answer this, we repeat the PCT score experiments with Llama3.2-1B (a smaller variant of the Llama3-8B model used before) in both quantized and non-quantized forms. Tables 13, 14, 15, 16, 17 (A.5) present the results. Overall trends hold across sizes and quantization: decoding parameters have minimal impacts on PCT scores, fine-tuning leads to significant shifts (as measured by t-tests), and the effects of prompt and fine-tuning (and their interactions) are substantial. However, in contrast to Llama3-8B-quantized, the prompt variation does not significantly affect Economic scores in Llama3.2-1B-quantized. Otherwise, the results are consistent across different model sizes and quantized and non-quantized versions of the same size, supporting generalizability.